what would the speed of an observer be if a red(4.688x10^14 Hz) traffic light appeared green (5.5555x10^14 Hz) to the observer ?
1
Expert's answer
2020-12-23T07:35:25-0500

Let's apply the expression for the relativistic transverse Doppler effect:


"\dfrac{f_t}{f_o}= \sqrt{\dfrac{1+\beta}{1-\beta}}"

where "f_t = 4.688\times10^{14} Hz" is the frequncy of the traffic light, "f_o = 5.5555\times10^{14} Hz" is the frequncy than the observer sees, "\beta = v/c", where "v" is the relative speed, and "c = 3\times 10^8m/s" is the speed of light in vacuum.

Expressing the "v" from the formula above, obtain:


"v = c\dfrac{f_t^2 - f_o^2}{f_t^2 + f_o^2}\\
v = 3\times 10^8\dfrac{(4.688\times10^{14})^2 - (5.5555\times10^{14})^2}{(4.688\times10^{14})^2 + (5.5555\times10^{14})^2}\approx \\
\approx -0.5\times 10^8m/s"

Answer. -0.5*10^8 m/s.